# Monocular Lane Detection Based on Deep Learning: A Survey

### v2

@@ Abstract
Lane detection plays an important role in autonomous driving perception systems. As deep learning algorithms gain popularity, monocular lane detection methods based on deep learning have demonstrated superior performance and emerged as a key research direction in autonomous driving perception. The core design of these algorithmic frameworks can be summarized as follows: (1) Task paradigm, focusing on lane instance-level discrimination; (2) Lane modeling, representing lanes as a set of learnable parameters in the neural network; (3) Global context supplementation, enhancing the detection of obscure lanes; (4) Perspective effect elimination, providing 3D lanes usable for downstream applications. From these perspectives, this paper presents a comprehensive overview of existing methods, encompassing both the increasingly mature 2D lane detection approaches and the developing 3D lane detection works. For a relatively fair comparison, in addition to comparing the performance of mainstream methods on different benchmarks, their inference speed is also investigated under a unified setting. Moreover, we present some extended works on lane detection, including multi-task perception, video lane detection, online high-definition map construction, and lane topology reasoning, to offer readers a comprehensive roadmap for the evolution of lane detection. Finally, we point out some potential future research directions in this field. We exhaustively collect the papers and codes of existing works at https://github.com/Core9724/Awesome-Lane-Detection and will keep tracing the research.
@@ I Introduction
Lane detection seeks to obtain the semantic and positional information of each lane line from the front view (FV) image captured by the onboard monocular camera. It is an indispensable part of the perception module in autonomous driving system, providing necessary prerequisites for subsequent decision-making and planning processes. Early lane detection methods depend on manually crafted operators for feature extraction [1, 2, 3, 4, 5, 6], which have limitations in accuracy and robustness when faced with complex scenes [7, 8, 9, 10, 11, 12, 13, 14, 15]. Later on, deep learning-based methods gradually dominate this field due to their strong feature representation ability and superior performance. Nowadays, deep learning-based monocular lane detection becomes a key research topic in autonomous driving perception, attracting great attention from both academia and industry.
@@
Figure 1: Two different technical routes to achieve complete lane detection. Different lane instances are visualized by different colors. The 2D lane detection results need to be combined with camera parameters after which the conversion to 3D lanes is done with the help of IPM. 3D lane detection methods directly incorporate the camera parameters into the network and achieve end-to-end prediction of 3D lanes.
@@
Figure 2: Structure of the overall review. Different colors represent specific sections. The summary of existing methods covers the core designs of the four lane detection algorithms mentioned in the introduction.
@@
Existing deep learning-based monocular lane detection methods can be divided into 2D lane detection and 3D lane detection methods. As shown in Figure 1, a complete lane detection process can be described as: given a FV image, the ultimate goal is to obtain 3D lane information in the ego vehicle coordinate system, i.e., the bird’s eye view (BEV) space. Because of the inherent perspective distortion in the camera imaging process, parallel lanes in the BEV plane intersect in the FV image, making it challenging to restore their true geometry. An effective solution involves building a 2D lane detection network [16, 17, 18, 19, 20] to obtain the 2D lanes from the FV image, after which the intrinsic and extrinsic parameters of the camera are combined to project these 2D lanes onto the ground through inverse perspective mapping (IPM) [21], thereby obtaining the final 3D lanes. With the rapid progress of deep learning-based methods, 2D lane detection achieve impressive results. However, IPM has an assumption that the ground is always flat, and does not consider the situations such as uphill, downhill, and rough road surfaces. In these cases, it leads to inaccurate 3D lane results, even if the 2D lane detection in FV is correct. Therefore, researchers are gradually shifting their focus to designing 3D lane detection networks [22, 23, 24, 25, 26, 27], which directly predict 3D lanes using FV images as input.
@@
For stable driving of vehicles in complex road environments, lane detection requires both precise localization and real-time processing capabilities. Apart from the above, the lanes which may be obscure due to nearby vehicles or severe weather conditions need to be detected. In order to better connect downstream tasks such as planning and control, the network also needs to differentiate between various lane instances and present the lane detection results in a vectorized format, such as an ordered set of points or a curve equation. This is because downstream requires calculating the driving planning line, i.e. the centerline, based on the lanes on both sides of the vehicle. It is difficult to perform calculations without distinguishing between different lane instances or lacking a vectorized representation format.
@@
Based on the complete lane detection process and the unique task requirements, the core design of the lane detection algorithmic frameworks can be summarized as follows: (1) Task paradigm, focusing on lane instance-level discrimination; (2) Lane modeling, representing lanes as a set of learnable parameters in the neural network; (3) Global information supplementation, enhancing the detection of obscure lanes; (4) Perspective effect elimination, providing 3D lanes usable for downstream applications. A comprehensive survey on lane detection should systematically explore these four key perspectives while integrating insights from both 2D and 3D lane detection methods. This dual focus ensures a holistic understanding of the advancements in lane detection technology, bridging the conceptual design with practical applications.
@@
$\bullet$ Related surveys. Early reviews on monocular lane detection mainly focused on traditional methods [28, 29]. Despite the existence of relevant summaries based on deep learning, the related survey papers [30, 31, 32] exhibited relatively narrow focus. On the one hand, only 2D lane detection methods or 3D lane detection methods are summarized. The close connection between 2D and 3D lane detection is ignored. On the other hand, the network structures or loss functions are paid too much attention. These aspects are also significant in deep learning, but they do not completely constitute the crux of addressing the lane detection challenge.
@@
$\bullet$ Contribution. To provide readers better understanding, we conduct a comprehensive investigation into the latest developments in monocular lane detection methods based on deep learning, starting from the core design of the lane detection algorithm frameworks. Compared to the related surveys [30, 31, 32], ours not only encompasses state-of-the-art 2D and 3D lane detection methods but also summarizes at a higher level. The main contributions of this survey can be followed as:
@@
    We present a comprehensive survey of deep learning-based monocular lane detection methods. This is the first survey that covers 2D lane detection and 3D lane detection.
@@
    This survey covers the core designs of four-lane detection algorithms: task paradigm (Distinguishing different lane instances), lane modeling (representing lanes as network learnable parameters), global information supplementation (identifying obscure lanes), and eliminating perspective effects (obtaining available 3D lanes for downstream). In this way, each type of method can be summarized into a universal pipeline.
@@
    In addition to reporting the performance indicators of representative methods, we also reevaluate the efficiency in a unified environment. This can enable readers to better compare various methods and help them choose appropriate baselines for their applications.
@@
$\bullet$ Organization. The rest of the survey is organized as Figure 2. Section II explains the datasets and evaluation metrics commonly used for Lane detection algorithms. Section III and Section IV review the existing 2D and 3D lane detection methods, respectively. We summarize existing methods from the perspective of the core design of lane detection algorithms. Section V reports on the performance of representative methods on typical datasets and efficiency comparisons in a unified environment, and analyzes them in conjunction with the core design of lane detection algorithms. We introduce some expanded works of lane detection in Section VI. Finally, in Section VII, we discuss possible future challenges for lane detection and provide conclusions in Section VIII.
@@ II-A Datasets
TABLE I: Monocular lane detection datasets at a glance. Under Region, “AS” denotes Asia, “NA” denotes North America, “Sim” denotes simulation data. Under Data Size, “Frames” denotes the number of annotated and total images, “Avg. Length” denotes the average time duration of videos. Under Diversity, “Inst. Anno.” denotes whether lanes are annotated at instance-level, “L. Cls.” denotes the category count (e.g. solid, dashed, etc.) of annotated lanes, and “Max L.” denotes the maximum number of lanes labeled in an image.
@@ II-A1 Datasets Commonly Used for 2D Lane Detection
$\bullet$ Tusimple. The TuSimple [35] dataset is collected with stable lighting conditions in highways, including different levels of occlusion, different types of lanes, and different road conditions. It only consists of 6,408 images, which are split into 3,268 training, 358 validation, and 2,782 test images. For each image, lanes are annotated by the 2D coordinates of sampling points with a uniform height interval of 10 pixels. Each annotated image has a size of 1280$\times$720 pixels.
@@ II-B Evaluation Metrics
As 2D lane detection benchmark representatives, Tusimple [35] and CULane [36] adopt two different ways to determine TP. Tusimple [35] focuses on point-by-point evaluation. The predicted point is considered correct if the horizontal distance from the true value point is less than 20 pixels when the longitudinal coordinates are the same. Furthermore, the line prediction is viewed as TP when it contains no less than 85% of the true value points. In contrast, CULane [36] emphasizes line-by-line evaluation, treating each lane as a mask of several pixels wide, and calculating the intersection (IoU) between the predicted lane and the annotated lane. The prediction with IoU larger than 75% is viewed as TP.
For 3D lane detection, there are also two main ways to determine the TP, represented by the evaluation methods in OpenLane [24] and ONCE-3DLanes [44] respectively. OpenLane [24] follow the evaluation metric designed by [23], the matching between prediction and ground truth is built upon edit distance, where one predicted lane is considered to be a TP only if 75% of its covered y-positions have a point-wise distance less than the max-allowed distance (1.5m). The ONCE-3DLanes [44] dataset employs a two-stage evaluation metric for lane detection. Initially, the IoU method from CULane [36] is utilized on the z-x plane (i.e., top view) to assess the alignment between the prediction and ground truth. Subsequently, if the IoU exceeds a predefined threshold, the curve matching error in camera coordinates is computed using unilateral chamfer distance. If this unilateral chamfer distance falls below the specified threshold, the prediction is classified as TP.
@@ III-A Classification Framework
Previous summaries [30, 31] primarily focus on the design of network structures and loss functions. It is noteworthy that instance-level discrimination and vectorized result representation, as unique requirements for the lane detection task and necessary prerequisites for guiding downstream applications, have been overlooked. By contrast, our classification of 2D lane detection methods is primarily according to the above two aspects.
As shown in Figure 3, first, for lane instance-level discrimination, 2D lane detection methods can be divided into two types of paradigms based on the number of stages required to complete the task: (a) Segmentation-based methods (two-stage), which complete lane recognition and instance discrimination in a certain order. Figure 4 summarizes the general pipeline for instance discrimination in such methods. (b) Object detection-based methods (one-stage), which can directly perform instance discrimination and localization concurrently. This advantage arises from the general pipeline of object detection algorithms, which execute both classification and regression tasks on a set of candidate proposals in parallel.
Second, vectorized result representation requires algorithms to consider how to model lanes as a set of values for neural network learning, i.e., lane modeling. In terms of lane modeling, the segmentation-based methods can be further divided into mask-based modeling, grids-based modeling, and keypoints-based modeling. For object detection-based methods, adopting a bounding box (bbox) to model a narrow and long lane is often not reasonable. This is because the bbox generated by object detection methods are mutually occluded, and there may be multiple instances of the same type of lane in the bbox. To align with the general object detection paradigm, these methods design unique ”bbox” to model lanes, including line anchor-based modeling and curve-based modeling. The details of each lane modeling are described in Figure 5.
Figure 3: The general pipelines for 2D lane detection: (a) Segmentation-based methods, which leverage mask, grids, or keypoints to model lanes. Each method only performs instance discrimination once. (b) Object detection-based methods, which leverage line anchor or parameter curve to model lanes and reach instance discrimination and localization in parall
Furthermore, most existing 2D lane detection datasets provide complete annotations for lanes, even in cases where the lanes are severely occluded by vehicles or affected by extreme weather conditions. This prompts many algorithms to meticulously design special structures within their networks to address this challenge, regardless of which classification system they belong to. Therefore, while reviewing these methods, the measures which they have taken to tackle it are also elaborated. We elaborately compare the representative 2D lane detection works according to the above classification criteria in Table II.
@@ III-B1 Mask-based modeling
Given an image $I\in R^{H\times W\times 3}$, the ultimate goal of the network is to predict a set of masks of the same size as the input image. In the early stages, fully convolutional segmentation networks represented by [49] are used to segment lanes [50]. The encoder extracts high-level semantic information into feature maps, and then the decoder upsamples these feature maps to their original size for pixel-wise prediction. However, even with more powerful general segmentation networks [51, 52, 53, 54, 55], lane segmentation performance remains unsatisfactory. This is largely due to these networks do not account for the elongated characteristics of lanes. Moreover, when lanes are occluded or interrupted by factors like lighting, relying solely on ground truth annotations of complete lanes is not an effective solution. Traditional encoders often fail to capture these subtle features. Consequently, numerous studies introduce specialized structures before pixel-wise prediction to enhance feature representation.
@@
VPGNet [34] predicts the disappearance points of lanes as the global geometric background to improve the performance of lane detection. SCNN [36] develops a novel convolutional layer for specially shaped objects, such as lanes and utility poles, allowing information to pass between image layers similarly to recurrent neural networks. However, there is room for improvement in computational speed. SAD [56] employs a self-attention distillation mechanism that contextually aggregates high-level and low-level attention to obtain finer lane features. IntRA-KD [57] uses a teacher-student distillation mechanism to represent lane structure knowledge as an interregional affinity map, capturing the similarity of lane feature distribution across different scene regions. EL-GAN [58] uses the generative adversarial network (GAN) to obtain more realistic and structurally rich lane segmentation results. Zhang et al. [59] combine feature fusion and multi-target segmentation-based on EL-GAN to design Ripple-GAN, which performs well for occluded lanes. Xu et al. [60] design a channel attention module that enhances lane features and suppresses background noise, and proposes a pyramid deformation convolution module to obtain more structural information of lanes. RESA [16] further proposes a recurrent aggregator on top of SCNN [36] that fully exploits the lane shape prior to enable the network to aggregate global features for improved performance and efficiency. PriorLane [61] obtains more comprehensive features based on a Mixed Transformer [62] and improves network performance by fusing image features and low-cost local prior knowledge, enhancing lane segmentation.
@@
While semantic segmentation provides semantic categories at the pixel level, it is insufficient for distinguishing different instances within the same category. An intuitive approach is to apply top-down instance segmentation frameworks, such as Mask R-CNN [63] or YOLACT [64], to achieve instance-level discrimination and segmentation of lanes. However, the bounding boxes generated by object detection methods may contain multiple lane instances, which complicates distinguishing them in the subsequent semantic segmentation process.
@@
SCNN [36] proposes a top-down manner. The approach is treating each lane in the image as a separate category and performing multi-category semantic segmentation. It incorporates a parallel classification branch to predict the existence of lanes at each position, subsequently combining the classification and segmentation results during post-processing to obtain the final lane. The subsequent work [58, 56, 16, 61] followed this method. However, these methods facilitate instance differentiation but introduce certain limitations: it requires defining a maximum number of lanes in advance to determine the number of possible instances. Additionally, the correspondence between lanes and classes is established by annotations. When vehicles switch between lanes, this predefined labeling may lead to ambiguity.
@@
To solve the above problem, some studies adopt a bottom-up approach for a more flexible instance segmentation, e.g., cluster the binary segmentation results of lanes/backgrounds. VPGNet [34] clusters lanes using a modified density-based clustering method. LaneNet [65] utilizes instance embedding to cluster the results of semantic segmentation, achieving lane instance segmentation. This method offers high clustering accuracy but is time-consuming, which limits its applicability for real-time vehicle experiments. FastDraw [70] constructs a learnable decoder that not only segments lanes but also identifies pixels belonging to the same lane. To address the inefficiency of pixel-embedding clustering, LaneAF [66] introduces an affinity vector field to associate pixels belonging to the same lane. Although these methods improve instance discrimination, the algorithm execution efficiency remains suboptimal due to the high complexity of bottom-up clustering and the low efficiency of mask-based modeling in classifying all pixels.
The lane mask obtained from network segmentation usually contains a large number of irrelevant areas. In order to be used for ego-vehicle motion prediction and planning, it is necessary to further denoise the mask to obtain vectorized results. Usually, for each lane mask, the highest response is sampled sequentially at equidistant heights, and then curve fitting is performed.

### v3

@@ Abstract
Lane detection plays an important role in autonomous driving perception systems. As deep learning algorithms gain popularity, monocular lane detection methods based on them have demonstrated superior performance and emerged as a key research direction in autonomous driving perception. The core designs of these algorithmic frameworks can be summarized as follows: (1) Task paradigm, focusing on lane instance-level discrimination; (2) Lane modeling, representing lanes as a set of learnable parameters in the neural network; (3) Global context supplementation, enhancing inference on the obscure lanes; (4) Perspective effect elimination, providing accurate 3D lanes for downstream applications. From these perspectives, this paper presents a comprehensive overview of existing methods, encompassing both the increasingly mature 2D lane detection approaches and the developing 3D lane detection works. Besides, this paper compares the performance of mainstream methods on different benchmarks and investigates their inference speed under a unified setting for fair comparison. Moreover, we present some extended works on lane detection, including multi-task perception, video lane detection, online high-definition map construction, and lane topology reasoning, to offer readers a comprehensive roadmap for the evolution of lane detection. Finally, we point out some potential future research directions in this field. We exhaustively collect the papers and codes of existing works at https://github.com/Core9724/Awesome-Lane-Detection and will keep tracing the research.
@@ I Introduction
Lane detection seeks to obtain the semantic and positional information of each lane line from the front view (FV) image captured by the onboard monocular camera. It is an indispensable part of the perception module in autonomous driving system, providing necessary prerequisites for subsequent decision-making and planning processes. Early lane detection methods depend on manually crafted operators for feature extraction [1, 2, 3, 4, 5, 6], which have limitations in accuracy and robustness when facing the complex scenes [7, 8, 9, 10, 11, 12, 13, 14, 15]. Later on, deep learning-based methods gradually dominate this field due to their strong feature representation ability and superior performance. Nowadays, deep learning-based monocular lane detection becomes a key research topic in autonomous driving perception, attracting great attention from both academia and industry.
@@
Figure 1: Two different technical routes to achieve complete lane detection. Different lane instances are visualized by different colors. The 2D lane detection results need to be projected to 3D space with the help of camera parameters and inverse perspective mapping (IPM). 3D lane detection methods directly incorporate the camera parameters into the network and achieve end-to-end prediction of 3D lanes.
@@
Figure 2: The structure of this paper. Different colors represent specific sections. The summary of existing methods covers four core designs of the lane detection algorithms mentioned in the introduction.
@@
Existing deep learning-based monocular lane detection methods can be divided into 2D lane detection and 3D lane detection methods. As shown in Figure 1, a complete lane detection process can be described as: given a FV image, the ultimate goal is to obtain 3D lane information in the ego vehicle coordinate system, i.e., the bird’s eye view (BEV) space. Because of the inherent perspective distortion in the camera imaging process, parallel lanes in the BEV plane intersect in the FV image, making it challenging to restore their true geometry. An effective solution involves building a 2D lane detection network [16, 17, 18, 19, 20] to obtain the 2D lanes from the FV image, then the intrinsic and extrinsic parameters of the camera are combined to project these 2D lanes onto the ground through inverse perspective mapping (IPM) [21], obtaining the final 3D lanes. With the rapid progress of deep learning-based methods, 2D lane detection achieve impressive results. However, IPM assumes that the ground is always flat and does not account for conditions like uphill, downhill, and rough road surfaces. It means that the IPM projection will lead the wrong 3D lane results, even if the 2D lane detection in FV is accurate. Therefore, researchers gradually shift their focus on designing 3D lane detection networks [22, 23, 24, 25, 26, 27], which directly predict 3D lanes using FV images as input.
@@
Precise localization and real-time processing are essential for lane detection. Apart from the above, autonomous vehicles must adapt to complex road environments where lanes may be obscure due to occlusion by nearby vehicles or adverse weather conditions. Lastly, to better connect downstream tasks like planning and control, each lane instance should be distinguished and presented in a vectorized format, such as an ordered set of points or a curve equation. This is because downstream requires calculating the driving planning lines, i.e., the centerlines, based on the lanes around the vehicle. It is difficult to perform calculations without distinguishing different lane instances or lacking a vectorized representation.
@@
Based on the complete process and the challenges of lane detection, the core design of the lane detection algorithmic frameworks can be summarized as follows: (1) Task paradigm, focusing on lane instance-level discrimination; (2) Lane modeling, representing lanes as a set of learnable parameters in the neural network; (3) Global context supplementation, enhancing the inference on the obscure lanes; (4) Perspective effect elimination, providing accurate 3D lanes for downstream applications. A comprehensive survey on lane detection should systematically explore these four key perspectives while integrating insights from both 2D and 3D lane detection methods. This dual focus is necessary to provide readers with a comprehensive understanding of advancements in lane detection technology, and help them to bridge the gap between conceptual design and practical applications.
@@
$\bullet$ Related surveys. Early reviews on monocular lane detection mainly focus on traditional methods [28, 29]. Despite the existence of relevant summaries based on deep learning, the related surveys [30, 31, 32] exhibit relatively narrow focus. On the one hand, only 2D or 3D lane detection methods are summarized. The close connection between 2D and 3D lane detection is ignored. On the other hand, the network structures or loss functions are paid too much attention in these papers, which are significant in deep learning but not the crux of addressing the lane detection challenges.
@@
$\bullet$ Contribution. This paper conducts a comprehensive investigation into the latest developments in monocular lane detection methods based on deep learning, focusing on the core designs of the lane detection algorithm frameworks. Compared to the related surveys [30, 31, 32], ours not only covers state-of-the-art 2D and 3D lane detection methods but also provides a higher-level summary. The main contributions of this survey can be summarized as follows::
@@
    We present a comprehensive survey of deep learning-based monocular lane detection methods. This is the first survey that covers both 2D lane detection and 3D lane detection.
@@
    This survey firstly introduces the four core designs of lane detection algorithms: task paradigm (Distinguishing different lane instances), lane modeling (representing lanes as network learnable parameters), global information supplementation (identifying obscure lanes), and eliminating perspective effects (obtaining available 3D lanes for downstream). In this way, each type of method can be summarized into a universal pipeline.
@@
    In addition to reporting the performance of representative methods, we also reevaluate their efficiency in a unified environment. This enables readers to more easily compare different methods and select the most suitable baselines for their applications.
@@
$\bullet$ Organization. The rest of the survey is organized as follows: Section II explains the datasets and evaluation metrics commonly used for Lane detection algorithms. Section III and Section IV reviews the existing 2D and 3D lane detection methods, respectively, where we summarize existing methods from the perspective of the core designs in lane detection algorithms. Section V reports on the performance of representative methods on typical datasets and efficiency comparisons in a unified environment, and analyzes them in conjunction with the core designs of lane detection algorithms. Some expanded works of lane detection are introduced in Section VI. The possible future challenges are discussed in Section VII and the conclusions are provided in Section VIII. The sturcture of this paper is shown in Figure 2.
@@ II-A Datasets
TABLE I: The overview of monocular lane detection datasets. For Region, “AS” denotes Asia, “NA” denotes North America, “Sim” denotes simulation data. For Data Size, “Frames” denotes the number of annotated and total images, “Avg. Length” denotes the average time duration of videos. For Diversity, “Inst. Anno.” denotes whether lanes are annotated at instance-level, “L. Cls.” denotes the category count (e.g. solid, dashed, etc.) of annotated lanes, and “Max L.” denotes the maximum number of lanes labeled in an image.
@@ II-A1 Datasets Commonly Used for 2D Lane Detection
$\bullet$ Tusimple. The TuSimple [35] dataset is collected with stable lighting conditions in highways, including different levels of occlusion, different types of lanes, and different road conditions. It consists of 6,408 images, which are split into 3,268 training, 358 validation, and 2,782 test images. For each image, lanes are annotated by the 2D coordinates of sampling points with a uniform height interval of 10 pixels. Each annotated image has a size of 1280$\times$720 pixels.
@@ II-B Evaluation Metrics
Tusimple [35] and CULane [36], which are the representive 2D lane detection benchmarks, adopts two different ways to determine TP. Tusimple [35] focuses on point-by-point evaluation. The predicted point is considered correct if the horizontal distance from the true value point is less than 20 pixels when the longitudinal coordinates are the same. Furthermore, the line prediction is viewed as TP when it contains no less than 85% of the true value points. In contrast, CULane [36] emphasizes line-by-line evaluation, treating each lane as a mask of several pixels wide, and calculating the intersection (IoU) between the predicted lane and the annotated lane. The prediction with IoU larger than 75% is viewed as TP.
For 3D lane detection, there are also two main ways to determine the TP, represented by the evaluation methods in OpenLane [24] and ONCE-3DLanes [44], respectively. OpenLane [24] follows the evaluation metric designed by [23]. The matching between prediction and ground truth is built upon edit distance, where one predicted lane is considered to be a TP only if 75% of its covered y-positions have a point-wise distance less than the max-allowed distance (1.5m). The ONCE-3DLanes [44] dataset employs a two-stage evaluation metric for lane detection. First, the IoU method from CULane [36] is utilized on the z-x plane (i.e., top view) to assess the alignment between the prediction and ground truth. Second, if the IoU exceeds a predefined threshold, the curve matching error in camera coordinates is computed using unilateral chamfer distance. If this unilateral chamfer distance falls below the specified threshold, the prediction is classified as TP.
@@ III-A Classification Framework
Previous summaries [30, 31] primarily focus on the design of network structures and loss functions. It is noteworthy that the instance-level discrimination and vectorized result representation, which are necessary prerequisites for guiding downstream applications, are overlooked. By contrast, our classification of 2D lane detection methods is primarily according to the above two aspects.
As shown in Figure 3, first, for lane instance-level discrimination, 2D lane detection methods can be divided into two types of paradigms based on the number of stages required to complete the task: (a) Segmentation-based methods (two-stage), which complete the lane localization and instance discrimination in a certain order. Figure 4 summarizes the general pipelines for instance-level discrimination in such methods. (b) Object detection-based methods (one-stage), which perform instance discrimination and localization concurrently. This advantage arises from the general pipeline of object detection algorithms, which execute both classification and regression tasks on a set of candidate proposals in parallel.
Second, vectorized result representation requires algorithms to consider how to model lanes as a set of values for neural network learning, i.e., lane modeling. In terms of lane modeling, the segmentation-based methods can be further divided into mask-based modeling, grids-based modeling, and keypoints-based modeling. For object detection-based methods, adopting a bounding box to model a narrow and long lane is often not reasonable. This is because the bounding boxes generated by object detection methods may be mutually occluded, and a bounding box may contain multiple lane instances. To align with the general object detection paradigm, these methods design unique ”bounding boxes” to model lanes, including line anchor-based modeling and curve-based modeling. The details of each lane modeling are described in Figure 5.
Figure 3: The general pipelines for 2D lane detection: (a) Segmentation-based methods, which leverage mask, grids, or keypoints to model a lane, perform instance discrimination and lane localization sequentially. Each method only performs instance discrimination once. (b) Object detection-based methods, which leverage line anchor or parameter curve to model a lane, complete the instance discrimination and lane localization in parallel.
Furthermore, most existing 2D lane detection datasets provide complete annotations for lanes, even the lanes are severely occluded by vehicles or affected by extreme weather conditions. To better identify such obscure lanes, many algorithms meticulously design special structures within their networks, thus the measures they use are also described in this section. We elaborately compare the representative 2D lane detection works according to the above classification criteria in Table II.
@@ III-B1 Mask-based modeling
Given an image $I\in R^{H\times W\times 3}$, the ultimate goal of the network is to predict a set of masks of the same size as the input image. In the early stages, fully convolutional segmentation networks represented by [49] are used to segment lanes [50]. The encoder extracts high-level semantic information into feature maps, and then the decoder upsamples these feature maps to their original size for pixel-wise prediction. However, even with more powerful general segmentation networks [51, 52, 53, 54, 55], the lane segmentation performance remains unsatisfactory. This is mainly because these networks do not account for the elongated nature of lanes. Moreover, when lanes are occluded by factors like vehicle or lighting, relying solely on annotations of complete lanes for supervision is not an effective solution. Traditional encoders often fail to capture these subtle features. Consequently, numerous studies introduce specialized structures before pixel-wise prediction to enhance feature representation.
@@
VPGNet [34] predicts the disappearance points of lanes as the global geometric background to improve the performance of lane detection. SCNN [36] develops a novel convolutional layer for specially shaped objects, such as lanes and utility poles, allowing information to pass between image layers, which is similar to recurrent neural networks. However, there is still some room for improvement in computational speed. SAD [56] employs a self-attention distillation mechanism that contextually aggregates high-level and low-level attention to obtain finer lane features. IntRA-KD [57] uses a teacher-student distillation mechanism to represent lane structure knowledge as an interregional affinity map, capturing the similarity of lane feature distribution across different scene regions. EL-GAN [58] uses the generative adversarial network (GAN) to obtain more realistic and structurally rich lane segmentation results. Then, Zhang et al. [59] select a GAN with better performance [60] and modify its structure to extract subtle lane features. Xu et al. [61] design a channel attention module that enhances lane features and suppresses background noise, and propose a pyramid deformation convolution module to obtain more structural information of lanes. RESA [16] further proposes a recurrent aggregator on top of SCNN [36] that fully exploits the lane shape prior to enable the network to aggregate global features for improved performance and efficiency. PriorLane [62] obtains more comprehensive features based on a Mixed Transformer [63] and improves network performance by fusing image features with low-cost local prior knowledge, enhancing lane segmentation.
@@
While semantic segmentation provides semantic categories at the pixel level, it is insufficient for distinguishing different instances within the same category. An intuitive approach is to apply top-down instance segmentation frameworks, such as Mask R-CNN [64] or YOLACT [65], to achieve instance-level discrimination and segmentation of lanes. However, the bounding boxes generated by object detection methods may contain multiple lane instances, which complicates distinguishing them in the subsequent semantic segmentation process.t
@@
SCNN [36] proposes a top-down process that is different from the above. Specifically, each lane is treated as a separate category so that multi-category semantic segmentation is performed. Meanwhile, a parallel classification branch is incorporated to predict the existence of lanes at each position. Finally, the classification and segmentation results are combined to obtain the final lanes. The subsequent works [58, 56, 16, 62] follow this way. This manner facilitates instance differentiation but introduces certain limitations: it requires defining a maximum number of lanes in advance to determine the number of possible instances. Additionally, the correspondence between lanes and classes is established by annotations. When vehicles switch between lanes, this predefined labeling may lead to ambiguity.
@@
To solve the above problem, some studies adopt a bottom-up approach for instance segmentation, i.e., cluster the binary segmentation results of lanes/backgrounds. VPGNet [34] clusters lanes using a modified density-based clustering method. LaneNet [66] utilizes instance embedding to cluster the results of semantic segmentation, achieving lane instance segmentation. This method offers high clustering accuracy but is time-consuming, which limits its applicability for real-time processing. FastDraw [71] constructs a learnable decoder that not only segments lanes but also identifies pixels belonging to the same lane. To address the inefficiency of pixel-embedding clustering, LaneAF [67] introduces an affinity vector field to associate pixels belonging to the same lane. Although these methods are more flexible, the algorithm execution efficiency remains suboptimal due to the high complexity of bottom-up clustering and the low efficiency of mask-based modeling in classifying all pixels.
The lane masks obtained from the segmentation network usually contain a large number of irrelevant areas. In order to be used for ego-vehicle motion prediction and planning, it is necessary to further denoise the mask to obtain vectorized results. Usually, for each lane mask, the highest response is sampled sequentially at equidistant heights, and then curve fitting is performed.